OpenStudy (ineedhelp27):

How many solutions are there to the equation below? 7(x+2)=7x-10

OpenStudy (ryanwhite499):

What do you think @INeedHelp27

OpenStudy (ryanwhite499):

Let's solve your equation step-by-step. 7(x+2)=7x−10 Step 1: Simplify both sides of the equation. 7(x+2)=7x−10 Simplify: (Show steps) 7x+14=7x−10 Step 2: Subtract 7x from both sides. 7x+14−7x=7x−10−7x 14=−10 Step 3: Subtract 14 from both sides. 14−14=−10−14 0=−24 Answer: There are no solutions.

OpenStudy (ryanwhite499):

So whats the answer?

OpenStudy (ryanwhite499):

If there are no solutions, and you ending up with a last part or step of the equation, what can be your answer

OpenStudy (ineedhelp27):

There was No Solution Thanks😄 A+

OpenStudy (ryanwhite499):

Welcome